Calculate the moment of inertia I of the following system of two masses connected by a light rod of length 4 meters rotating about an axis O, 1 meter from mass 3kg and 3 meters from mass 5 kg.
1
Expert's answer
2012-11-27T07:25:27-0500
I = m1*r1^2 + m2*r2^2 = (3kg)*(1 m)^2 + (5 kg)*(3 m)^2 = 48 kg*m^2.
The moment of inertia of the rod is not accounted for, as we're assuming its
mass to be zero.